# Building Better: Avoiding Pitfalls in Developing Language Resources when Data is Scarce

### v5

@@ 4.3 Credit Attribution
We asked the respondents to share whether they were properly credited for their work by, e.g., getting financial compensation for a long annotation task, getting involved in the writing of a research paper for a resource that they built, etc. As shown in Figure 5, most respondents (>67%) report this not being the case at least once. Figure 6 shows the distributions of responses pertaining to how the respondents were incentivised to perform an annotation task for which they were eventually not given due credit.
@@
For the respondents who reported that they did not receive proper credit for projects in which they were involved, we report the incentives for joining these projects and the time it took the participants to complete the work. As shown in Figure 6, they were either:
@@
The Issue with the Over-reliance on Online Communities
When using standard crowdsourcing platforms such as AMT or Prolific, one can operationalise the annotation for a given task. Despite their shortcomings Fort et al. (2011); Irani (2015), one can attempt to protect workers by using tests and training when annotating hard tasks. However, for mid- to low-resource languages, platforms such as AMT and Prolific often do not have enough speakers registered on the platform. Therefore, researchers opt for personal connections or participatory efforts instead. On the one hand, personal outreach and community efforts make people feel more included and trust can be built more easily. On the other hand, there is a high risk of exploitation and emotional manipulation in such a case. For instance, some junior researchers reported cases of exploitation due to them being told that joining an online community that builds language resources is prestigious and worth adding to their CVs. We note that some respondents shared their frustration in the responses as they were offered, e.g., small company gifts for months of labor. As shown in Figure 6, 40% of the respondents, who spent 1 day to more than a month annotating data report negative experiences. That is, their work was not properly compensated, acknowledged, or recognised. This calls for a need to set guidelines and standards when using community or participatory research services.
@@ 5 Recommendations
While there has been a considerable amount of work on the ethics of best practices for building NLP and ML artefacts Bender and Friedman (2018); Leech et al. (2024); Mohammad (2022, 2023), our findings substantiate the fact that research on under-resourced languages presents additional challenges linked to the reliance on unconventional practices.
@@ 5.1 Center the People
Our findings show that various issues ought to be addressed early as research in the area lacks established standards and is subject to power differentials (e.g., some researchers reported their work being used by more notable individuals for fame). Many under-resourced languages are from what is called “the Global South” with a large number of them being spoken rather than written.
@@
Language is an important part of a population’s identity and technologies dealing with it have a direct impact on people’s lives. Past NLP work highlights how to engage with speakers and communities whose languages are in question Bird (2020); Bird and Yibarbuk (2024); Bird (2024); Cooper et al. (2024); Ramponi (2024). We further reinforce this argument with our findings, i.e., most researchers researchers report marked limitations in their languages (see Figure 4). Hence, when a researcher reaches out to a group of people with little background knowledge of their culture or language, one needs to approach these problems from the perspective of the community in question Bird (2022). The question of who is exactly served needs to be addressed early on to avoid any misconception of perceived needs for language technologies.
@@
In addition to the large percentage of our survey respondents who reported not being properly credited for their labor (Figure 5), there were cases of emotional manipulation (e.g., making emotional arguments such as how one’s labor will help the speakers of the language and that is compensation enough). One has to set rules and expectations with clear communication on the purpose of a given research project. For instance, when dealing with online participatory research communities for data collection and annotation, extra care needs to be shown and benevolent prejudice such as depicting oneself as a savior of a local community Bird (2022) must be avoided. Companies and research labs relying on communities for annotation and data creation need to properly compensate the contributors by, e.g., drawing legal contracts. The question of who is annotating what has to be addressed as well. The scarcity of annotators can lead to poor choices as very often, native speakers cannot be found online easily which has led to researchers choosing people from associated regions—people who do not necessarily speak the language variant in question. This results in a problematic overgeneralisation that puts different languages under the same umbrella simply because they have one or a small set of attributes in common. This often results in potentially over-simplistic solutions. For instance, variations of Arabic differ considerably but numerous research projects have treated entire regions, such as North Africa, as a monolith (e.g., to appear to have more data).
@@ 5.2 Be Fair: Give Credit where Credit is Due
Our findings show the unfortunate trend of data workers and NLP researchers suffering from a lack of recognition, especially those who are part of online communities (Figure 3). A needed follow-up work would be extensive fieldwork with the various online communities. Hence, our recommendation is a call for action on the setup of fair and comprehensive practices in participatory research projects, while taking power differentials into account. That is, existing authorship standardsneed to be followed and discussed prior to the start of a project as to whether a data worker should be listed as an author. This is particularly critical for junior researchers who substantially contribute to the resource construction and may assume that they will have a leading role in the paper. Moreover, proper financial compensation needs to be provided for annotators. Ideally, a resource paper should provide proof that the annotators were paid and treated fairly as recommended by Rogers et al. (2021).
@@ 5.3 Choose the Jargon Carefully and Be Aware of False Generalisations
As previously discussed in 5.1, it is important to embrace social awareness and avoid grouping people from colonial and Western perspectives Bird (2020, 2022); Held et al. (2023). Hence, one can avoid dismissive and outdated terms and classifications, e.g., “the rest of the world”. Note also that The World’s Values Survey classification Haerpfer and Kizilova (2012), which is often used in NLP papers (e.g., Santy et al. (2023)), presents an orientalist view of the world Said (1977). It has clear flaws such as including Christian-majority countries (e.g., Ethiopia, Rwanda) in a so-called "African-Islamic" category and groups of countries that have little in common in one category (e.g., Kyrgyzstan and Tunisia), leading to misrepresentations and enforcing stereotyping.
@@ 5.4 Set Fair and Realistic Expectations
As pointed out by Doğruöz and Sitaram (2022), tools for low-resource languages are often perceived as scaled-down versions of high-resource ones. Adding to previous work elaborating on what this may mean to the speakers Bird (2022); Markl et al. (2024), we focus on the impact of setting these expectations for researchers and practitioners. That is, they may be expected to build models similar to those built for high-resource languages, i.e., tackling the same NLP tasks, and performing extremely well. However, this can be unrealistic for various reasons such as the users’ needs Blaschke et al. (2024), the language’s specific features Bird and Yibarbuk (2024), and the lack of funding linked to the “prestige” of the language as reported by our respondents and similarly discussed by Mihalcea et al. (2024) in the context of LLM research.
@@
Joshi et al. (2020) conducted a survey on the state of NLP in various languages. They found that people do not necessarily want the tools that researchers think they need. However, we should not be limiting what NLP research on low- to mid-resource languages should be about. The most challenging part is to find a balance between what technologies local communities need and the research explored by the NLP community. This balance could be reinforced by continued communication with local communities Bird and Yibarbuk (2024); Lent et al. (2022); Mager et al. (2023); Cooper et al. (2024).
Dealing with a "Solved" Problem in a New Language is an Actual Contribution
Just because automatic systems obtain high scores on some English datasets (often making use of massive amounts of English data) does not mean that the problem is solved generally, or that working on that problem in other languages is no longer interesting or valuable. Working with each new language presents many new challenges, e.g., a rich morphology or the presence of tones Adebara and Abdul-Mageed (2022). Therefore, such work should not be undervalued by calling it “a replication” study.
@@ 5.5 Check the Source Even if the Language is Low-resource
Due to the limited amount of online data available for some languages, there is a tendency to use any available online source to build resources for these languages without examining the ethical implications or the appropriateness of the source. While it is typically easier to use religious texts, song lyrics, or movie subtitles, these should be carefully considered Hutchinson (2024); Mager et al. (2023). For instance, song lyrics are not representative of daily communication Mayer et al. (2008) since, e.g., they often rhyme and contain profanities. Additionally, using religious texts without a thorough inspection of potential implications can lead to misrepresentations (e.g., considering some cultures to be highly religious) Mager et al. (2023). Further, some work turns to synthetic data generated using machine translation and LLMs when these show clear limitations, especially in multicultural settings Hershcovich et al. (2022)(see Section 4.2.1). Research from other disciplines, even tightly related, such as linguistics Turner (2023), can help us choose adequate and suitable data sources.
@@ 6 Conclusion
We present insights from NLP researchers and practitioners working on under-served languages. We discuss limitations and research practices in the field, then provide actionable recommendations on how to address the reported issues while remaining fair to data workers. Our work is the first to document NLP researchers and workers’ experiences.
@@
We acknowledge the fact that there are experiences that are different from those of our respondents and the risk of selection bias. Nonetheless, it is also important to give voice to the concerns of data annotators and researchers working on mid- to low-resource languages, and our survey and this paper aim to do that.
@@
While most respondents shared their contact information, it was mainly for following up on the resulting study. We do not share any information that may reveal their identity or the projects they reported.
@@ 7 Acknowledgments
We thank our survey respondents for providing insightful comments and we are grateful to anyone who helped us spread the word on social media or by email.
Many thanks to Joanne Boisson, Michael Schlichtkrull, Sowmya Vajjala, and Evani Radiya-Dixit for their helpful feedback.
Appendix A Appendix
A.1 Questionnaire
We would like to investigate the common practices in NLP research on low-resource languages (language variants and "dialects" included).
@@ A.1 Questionnaire
  • (Optional long text answer) Can you name the institutions/projects? (We will not make the names public if you do not want to share the names publicly. See question below.)

### v6

@@ 4.3 Credit Attribution
We asked the respondents to share whether they were properly credited for their work by, for instance, getting financial compensation for a long annotation task, getting involved in the writing of a research paper for a resource that they built, etc. As shown in Figure 5, most respondents (>67%) report this not being the case at least once. Figure 6 shows the distributions of responses pertaining to how the respondents were incentivised to perform an annotation task for which they were eventually not given due credit.
@@
For the respondents who reported that they did not receive proper credit for projects in which they were involved, we list the main incentives for joining these projects and the time it took the participants to complete the work. As shown in Figure 6, they were either:
@@
The Issue with Over-Reliance on Online Communities and Participatory Research
When using standard crowdsourcing platforms such as AMT or Prolific, researchers can operationalise the annotation process for a given task. Despite their shortcomings Fort et al. (2011); Irani (2015), it is possible to protect workers by implementing tests and training for complex annotation tasks. However, for mid- to low-resource languages, platforms such as AMT and Prolific often lack sufficient numbers of registered speakers. As a result, researchers frequently resort to personal networks or participatory approaches. On the one hand, personal connections and community engagement can foster a sense of inclusion and help build trust. On the other hand, however, these approaches carry a significant risk of exploitation and emotional manipulation. That is, some junior researchers were clearly exploited, having been previously told that joining an online community for building language resources was prestigious and worth adding to their CVs. Some respondents expressed frustration at being offered, for instance, company merchandise in exchange for months of labor—a situation that raises further concerns about the involvement of big tech in NLP research for low-resource languages Abdalla et al. (2023). As shown in Figure 6, 40% of respondents who spent between one day and over a month annotating data reported negative experiences. Their contributions were not adequately compensated, acknowledged, or recognised. This highlights the urgent need for clear guidelines and standards when engaging in community-based or participatory research efforts.
@@ 5 Recommendations
While there has been a considerable amount of work on the ethics of best practices for building NLP and ML artefacts Bender and Friedman (2018); Mohammad (2022, 2023); Leech et al. (2024), our findings substantiate the fact that research on under-resourced languages presents additional challenges linked to the reliance on unconventional practices.
@@ 5.1 Center the People
Our findings show that various issues ought to be addressed early as research in the area lacks established standards and is subject to power differentials. For example, some researchers reported that their work and contributions were diminished by more prominent individuals taking first authorship on papers to which they had contributed little or not at all. Many under-resourced languages are from what is called “the Global South” with a large number of them being spoken rather than written.
@@
Language is an important part of a population’s identity and technologies dealing with it have a direct impact on people’s lives. Past NLP work highlights how to engage with speakers and communities whose languages are in question Bird (2020); Bird and Yibarbuk (2024); Bird (2024); Cooper et al. (2024); Ramponi (2024). We further reinforce this argument with our findings, i.e., most respondents report marked limitations in their languages (see Figure 4). Hence, when a researcher reaches out to a group of people with little background knowledge of their culture or language, one needs to approach these problems from the perspective of the community in question Bird (2022). The question of who is exactly served needs to be addressed early to avoid any misconception of perceived needs for language technologies.
@@
In addition to the large proportion of survey respondents who reported not being properly credited for their labor (Figure 5), there were instances of emotional manipulation. This included appeals suggesting that one’s labor would benefit the speakers of the language, and that this alone should be considered sufficient compensation. Notably, such practices were common in participatory research projects initiated by both small groups and Big Tech companies. One has to set rules and expectations with clear communication on the purpose of a given research project. For instance, when dealing with online participatory research communities for data collection and annotation, extra care needs to be shown and benevolent prejudice such as depicting oneself as a savior of a local community Bird (2022) must be avoided.
The question of who is annotating what has to be addressed as well. The scarcity of qualified annotators can result in poor decision-making. Native speakers are often difficult to find online, which has led some researchers to select individuals from broadly associated regions—people who do not necessarily speak the specific language variant in question. This results in problematic overgeneralisations and overly simplistic solutions, where distinct languages or dialects are grouped together simply because they share one or a few attributes. For example, while Arabic dialects vary significantly, numerous research projects have treated entire regions, such as North Africa, as a linguistic monolith, sometimes to appear to have more data than is truly available.
@@ 5.2 Be Fair: Give Credit where Credit is Due
Our findings highlight an unfortunate trend—data workers and NLP researchers, especially those collaborating on participatory research projects, often suffer from a lack of recognition (Figure 3). Accordingly, our recommendation is to set fair and comprehensive practices in participatory research projects, while considering power differentials.
Monetary Compensation
Annotators must be provided with proper financial compensation. Companies and research labs that rely on communities for annotation and data creation should ensure fair compensation for contributors, for example, through legally binding contracts. Ideally, resource papers should include evidence that annotators were paid and treated fairly, as recommended by Rogers et al. (2021).
Co-authorship Standards
Existing authorship standards must be followed and discussed before the start of a project, particularly with respect to whether data workers should be listed as authors. This is especially important for junior researchers who contribute significantly to resource construction and may expect a leading role in the resulting publication. To inclusively recognise the contributions of junior researchers and those involved in data curation, we recommend the following:
  1. 1.
    Credit all contributors involved in dataset creation, including those that contributed to data curation, design of the annotation setup, and the writing of the paper—as authors.
  2. 2.
    Let data annotators know that they are welcome to play a greater role in the project, possibly rising to the level of co-authors. Discuss with them that meaningful contributions to one or more of the following can lead to co-authorship:
    • Participating in data curation and experimental design.
    • Running language-specific experiments and performing (language-specific) error analyses.
    • Conducting (language-specific) ablation studies.
    • Contributing to the writing of specific paper sections, such as the related work.
    • Providing insights into the resource by:
      • Selecting culturally relevant or representative examples.
      • Offering explanations and interpretations.
      • Describing the annotation process and sharing key observations or challenges.
These tips are adapted from large past annotations efforts that we led Ousidhoum et al. (2024); Muhammad et al. (2025).
@@ 5.3 Choose the Jargon Carefully and Be Aware of False Generalisations
As previously discussed in 5.1, it is important to embrace social awareness and avoid grouping people from colonial and Western perspectives Bird (2020, 2022); Held et al. (2023). Hence, one can avoid dismissive and outdated terms and classifications, e.g., “the rest of the world”. Note also that The World’s Values Survey classification Haerpfer and Kizilova (2012), which is often used in NLP papers (e.g., Santy et al. (2023)), presents an orientalist view of the world Said (1978). It has clear flaws such as including Christian-majority countries (e.g., Ethiopia, Rwanda) in a so-called “African-Islamic” category and groups of countries that have little in common in one category (e.g., Kyrgyzstan and Tunisia, Bolivia and the Philippines, South Africa and the UAE), leading to misrepresentations and enforcing stereotypes.
@@ 5.4 Set Fair and Realistic Expectations
As pointed out by Doğruöz and Sitaram (2022), tools for low-resource languages are often perceived as scaled-down versions of those developed for high-resource languages. Building on previous work that explores what this may mean for speakers Bird (2022); Markl et al. (2024), we shift the focus to researchers and practitioners, who are often expected to build similar models to those created for high-resource languages—i.e., tackling the same NLP tasks and achieving similarly high performance. However, this expectation can be unrealistic for several reasons, including users’ actual needs Blaschke et al. (2024), the specific linguistic features of the language Bird and Yibarbuk (2024), and the lack of funding often tied to the perceived “prestige” of a language, as reported by our respondents and similarly discussed by Mihalcea et al. (2024) in the context of LLM research.
@@
Joshi et al. (2020) conducted a survey on the state of NLP across various languages and found that people do not necessarily want the tools that researchers assume they need. One should not prescribe what NLP research on mid- to low-resource languages should be about. The real challenge lies in striking a balance between the technologies that local communities require and the directions pursued by the research community. This balance can be strengthened through ongoing communication and collaboration with various stakeholders Lent et al. (2022); Mager et al. (2023); Bird and Yibarbuk (2024); Cooper et al. (2024).
Dealing with a “Solved” Problem in a New Language is an Actual Contribution
Just because automatic systems obtain high scores on some English datasets, often making use of massive amounts of English data, does not mean that the problem is solved generally, or that working on that problem in other languages is no longer interesting or valuable. Working with each new language presents many new challenges, e.g., a rich morphology or the presence of tones Adebara and Abdul-Mageed (2022). Such work should not be undervalued by calling it “a replication” study.
@@ 5.5 Check the Source Even if the Language is Low-resource
Due to the limited availability of online data for some languages, there is a tendency to use any accessible source to build resources—often without considering the ethical implications or the appropriateness of the content. While it is typically more convenient to rely on religious texts, song lyrics, or film subtitles, such sources should be critically assessed Hutchinson (2024); Mager et al. (2023). For example, song lyrics are not representative of everyday communication Mayer et al. (2008), as they often rhyme and may contain profanity that is not typical of daily language use.
Moreover, scraping Indigenous texts without obtaining proper consent or permissions, and without sensitivity to cultural or religious contexts Hutchinson (2024), often results in work on low-resource languages being framed in utilitarian terms, with insufficient attention to deontological concerns. For instance, the use of religious texts without acknowledging their potential implications can lead to cultural misrepresentations, such as portraying certain communities as uniformly religious Mager et al. (2023).
Additionally, some researchers resort to synthetic data generated via machine translation or large language models (LLMs), despite well-documented limitations, particularly in multicultural or multilingual contexts Hershcovich et al. (2022) (see Section 4.2.1). Insights from other disciplines, including closely related fields such as linguistics Turner (2023), can help guide the selection of more culturally appropriate data sources, especially considering that such datasets are likely to persist and influence future research.
5.6 Position Your Contribution
Similar to Hutchinson (2024), we encourage the inclusion of positionality statements. Specifically, authors should indicate their relationship to the language(s) they work on. This may include, for example, their level of fluency, whether they have formally studied the language(s) in question, and whether they collaborated with native or fluent speakers for tasks such as data annotation. It is also important to clearly state the cultural background of data creators and annotators to avoid false generalisations and regional marginalisation (e.g., treating all Arabic-speaking countries as one homogeneous group Keleg (2025)).
@@ 6 Conclusion
We present insights from NLP researchers and practitioners working on under-served languages. We discuss reported limitations in this area of research and highlight issues related to data standards, along with common practices and concerning trends, such as problematic incentivisation and the lack of recognition for data workers’ labor.
We then offer actionable recommendations to improve data quality by reflecting on data sources, setting realistic expectations for under-resourced languages, and centering speakers and end users while ensuring fairness to data workers.
@@
We acknowledge the the risk of selection bias. Nonetheless, our main goal was to give voice to the concerns of data annotators and researchers working on mid- to low-resource languages.
@@
While most respondents shared their contact information, it was mainly for following up on the resulting study.
We do not share any information that may reveal their identities or the projects they reported on.
Positionality Statement
The three authors are affiliated with governmental and academic institutions in the UK, Sweden, and Canada, respectively. They have experience working with large, diverse groups and in developing new NLP artefacts for high-, mid-, and low-resource languages.
Nedjma Ousidhoum is linguistically proficient in Arabic (Algerian, MSA, and Classical), French, and English. She has lived in Algeria, Hong Kong, and the UK.
Meriem Beloucif is linguistically proficient in Arabic (Algerian and MSA), French, and English, and has a moderate knowledge of German and Swedish. She has lived in Algeria, Hong Kong, Denmark, Germany, and Sweden.
Saif M. Mohammad is proficient in English, Urdu, and Hindi. He has lived in India, USA, and Canada.
@@ 7 Acknowledgments
We thank our survey respondents for their insightful comments, and we are grateful to everyone who helped us spread the word on social media or via email.
Many thanks to Joanne Boisson, Michael Schlichtkrull, Sowmya Vajjala, and Evani Radiya-Dixit for their helpful feedback, as well as to the anonymous Action Editor and reviewers for their comments and suggestions.
Appendix
Appendix A Questionnaire
We would like to investigate the common practices in NLP research on low-resource languages (language variants and “dialects” included).
@@ Appendix A Questionnaire
  • (Optional long text answer) Can you name the institutions/projects? (We will not make the names public if you do not want to share them publicly. See question below.)